OpenStudy (blackasinc):

Find all values of x for which the following is true. sin x = 1 a. π/3 + 3k π b. -π/3 +2k π c. -π/2 +2k π d. π/2 + 2k π e. π/4 + 2k π How do i figure this out?

OpenStudy (legomyego180):

do you know how to find x here?

OpenStudy (blackasinc):

No. Im not sure.

OpenStudy (marcelie):

what x value gives u 1 on the unit circle

OpenStudy (legomyego180):

Firstly, you need to have the unit circle memorized to do this problem

OpenStudy (legomyego180):

nope thats cosine

OpenStudy (blackasinc):

π/4

OpenStudy (legomyego180):



OpenStudy (legomyego180):

Whats sin(pi/2)?

OpenStudy (blackasinc):

0.027

OpenStudy (blackasinc):

sin (90) = 1

OpenStudy (legomyego180):

That's in degrees, if you can use your calculator for this problem, make sure you have it set to radian mode by pressing [Mode] then setting it from degree to radians. So the reason you cant do this problem is because you havent memorized the unit circle. Memorize this and it will make your life 100x easier for the effort. You have to use it in every other math class you take from now on until forever so might as well get it over with. http://www.shelovesmath.com/wp-content/uploads/2012/11/Unit-Circle.png?36349e Sin of pi/2 = 1

OpenStudy (legomyego180):

So, which answer has pi/2 in it

OpenStudy (blackasinc):

D

OpenStudy (blackasinc):

ive been practicing memorizing th unit circle today, but was not sure how this problem related to it.

OpenStudy (sshayer):

\[\sin x=1,\sin x=1=\sin \frac{ \pi }{ 2 }=\sin \left( 2 n \pi +\frac{ \pi }{ 2 } \right)\] \[x= 2n \pi +\frac{ \pi }{ 2 },\] where n is an integer.

OpenStudy (legomyego180):

the 2n*pi is just used to account for all coterminal angles. The pi/2 is the important bit